OpenStudy (darkbluechocobo):

Calculate the time-of-flight for a projectile launched at an angle to the horizontal if the velocity is 141.4m/s and the angle is 45 degrees

OpenStudy (darkbluechocobo):

@amistre64

OpenStudy (anonymous):

The time of flight is when y position of object is equal to zero, therefor using the equation of motion: \[y=-\frac{ 1 }{ 2 }gt ^{2}+V _{o}y(t) \] and the given info: V = 141.4 m/s angle = 45 first we calculate the Y component of velocity: Vy = 141.4 * Sin(45) = 99.98 m/s then we substitute the Y component of velocity and y position (which is zero) and "g" (g = 9.8 m/s) in the above equation to get the following: \[0=-\frac{ 1 }{ 2 }(9.8)t ^{2}+99.98t \] not its just algebra, first divide the whole thing by "t", then take 99.98 to the other part, divide be "t" coefficient to get t the answer is: t = 20.4s

OpenStudy (darkbluechocobo):

Thank you :D